# Gradient-based Sample Selection for Faster Bayesian Optimization

### v1

@@ Abstract
Bayesian optimization (BO) is an effective technique for black-box optimization. However, its applicability is typically limited to moderate-budget problems due to the cubic complexity in computing the Gaussian process (GP) surrogate model. In large-budget scenarios, directly employing the standard GP model faces significant challenges in computational time and resource requirements. In this paper, we propose a novel approach, gradient-based sample selection Bayesian Optimization (GSSBO), to enhance the computational efficiency of BO. The GP model is constructed on a selected set of samples instead of the whole dataset. These samples are selected by leveraging gradient information to maintain diversity and representation. We provide a theoretical analysis of the gradient-based sample selection strategy and obtain explicit sublinear regret bounds for our proposed framework. Extensive experiments on synthetic and real-world tasks demonstrate that our approach significantly reduces the computational cost of GP fitting in BO while maintaining optimization performance comparable to baseline methods.
@@ 1 Introduction
Bayesian optimization (BO) (Frazier, 2018) is a successful approach to black-box optimization that has been applied in a wide range of applications, such as hyperparameter optimization, reinforcement learning, and mineral resource exploration. BO’s strength lies in its ability to represent the unknown objective function through a surrogate model and by optimizing an acquisition function (Garnett, 2023; Wang et al., 2023). BO consists of a surrogate model, which provides a global predictive model for the unknown objective function, and an acquisition function that serves as a criterion to strategically determine the next sample to evaluate. In particular, the Gaussian process (GP) model is often preferred as the surrogate model due to its versatility and reliable uncertainty estimation. However, the GP model often suffers from large data sets, making it more suitable for small-budget scenarios (Binois & Wycoff, 2022). To fit a GP model, the dominant complexity in computing the inversion of the covariance matrix inversion is $\mathcal{O}(n^{3})$, where $n$ is the number of data samples. As the sample set grows, the computational burden increases substantially. This limitation poses a significant challenge for scaling BO to real-world problems with large sample sets.
@@
Despite the various approaches in improving the computational efficiency of BO, including parallel BO (González et al., 2016; Daulton et al., 2021, 2020; Eriksson et al., 2019), kernel approximation methods (Kim et al., 2021; Jimenez & Katzfuss, 2023; Hensman et al., 2013; Williams & Seeger, 2000) and sparse GP (Lawrence et al., 2002; Leibfried et al., 2020; McIntire et al., 2016), the computational overhead can still become a burden in practice. Current implementations of sparse GPs (McIntire et al., 2016) for BO adopt iterative schemes that add a new sample while removing one from the original subset at each iteration. Although this approach reduces computational complexity, it can be suboptimal in identifying the most representative subset, especially in complex optimization landscapes. Furthermore, while BO algorithms are theoretically designed to balance exploitation and exploration, with limited budget in practice, they can over-exploit current best regions before shifting to exploration Wang & Ng (2020), leading to suboptimal performance in locating the global optimums.
During the iterative search process of BO, some samples can become redundant and contribute little to the additional information gain. Such samples collected in earlier stages thus diminish in importance as the process evolves. For instance, excessive searching around identified minima becomes redundant once the optimal value has been determined, as these samples cease to offer meaningful insights for further optimization. To efficiently fit a GP, it is essential to focus on samples that provide the most informative contributions. As shown in Figure 1, carefully selected samples can effectively fit a GP. Despite the reduced number of samples, the GP still captures the key trends and features of the true function while maintaining reasonable uncertainty bounds. In this paper, we propose to incorporate the gradient-based sample selection technique into the BO framework to enhance its scalability and effectiveness in large-budget scenarios. This technique was originally proposed for continuing learning with online data stream (Aljundi et al., 2019). The previously seen data are selectively sampled and stored in a replay buffer to prevent catastrophic forgetting and enhance model fitting. The iterative optimization process of BO can be viewed as an online data acquiring process. Hence, the GP model fitting in the next iteration can be seen as a similar continuing learning problem. By using gradient information to gauge the value of each sample, one can more judiciously decide which samples are most essential for building a subset, and maintain the most representative subset of BO samples. This subset is then used to fit the GP model, accelerating the BO process while ensuring efficient and effective GP fitting and mitigating the problem of over-exploitation. We summarize our main contributions as follows:
@@
  • Theoretical analysis. We provide a rigorous theoretical analysis of the regret bound for the GSSBO. Theoretical results show that the regret bound of the proposed algorithm (with sample selection) is similar to that of the standard GP-UCB algorithm (without sample selection).
  • Empirical validations. We conduct comprehensive numerical experiments, including synthetic and real-world datasets, to demonstrate that compared to baseline methods, the proposed algorithm achieves comparable performance, but significantly reduces computational costs. These results verify the benefit of using gradient information to select a representative subset of samples.
@@ 2 Related Works
BO with Resource Challenges. In practical applications, BO faces numerous challenges, including high evaluation costs, input-switching costs, resource constraints, and high-dimensional search spaces. Researchers have proposed a variety of methods to address these issues. For instance, parallel BO employs batch sampling to improve efficiency in large-scale or highly concurrent scenarios (González et al., 2016; Daulton et al., 2021, 2020; Eriksson et al., 2019). Kernel approximation methods, such as random Fourier features, map kernels onto lower-dimensional feature spaces, thus accelerating kernel-based approaches (Rahimi & Recht, 2007; Kim et al., 2021). Multi-fidelity BO leverages coarse simulations together with a limited number of high-fidelity evaluations to reduce the overall experimental cost (Kandasamy et al., 2016). For high-dimensional tasks, techniques such as random embeddings or active subspaces help reduce the search dimensionality (Wang et al., 2016). Meanwhile, sparse GP significantly reduces computational complexity by introducing “inducing points” (Lawrence et al., 2002; Leibfried et al., 2020; McIntire et al., 2016). However, these methods face limitations in practical usage scenarios, often struggling to balance complex resource constraints while dynamically adapting to high-dimensional and rapidly changing environments.
@@
BO with Gradient Information. The availability of derivative information can significantly simplify optimization problems. Ahmed et al. (2016) highlight the potential of incorporating gradient information into BO methods and advocate for its integration into optimization frameworks. Wu & Frazier (2016) introduced the parallel knowledge gradient method for batch BO, achieving faster convergence to global optima. Rana et al. (2017) incorporated GP priors to enable gradient-based local optimization. Chen et al. (2018) proposed a unified particle-optimization framework using Wasserstein gradient flows for scalable Bayesian sampling. Bilal et al. (2020) demonstrated that BO with gradient-boosted regression trees performed well in cloud configuration tasks. Tamiya & Yamasaki (2022) developed stochastic gradient line BO (SGLBO) for noise-robust quantum circuit optimization. Penubothula et al. (2021) funded local critical points by querying where the predicted gradient is zero. Zhang & Rodgers (2024) introduced BO of gradient trajectory (BOGAT) for efficient imaging optimization. Although these methods leverage gradient information to improve optimization efficiency and performance, they mainly focus on refining the GP model or acquisition functions.
@@
Subset Selection. Subset selection is a key task in fields such as regression, classification, and model selection, aiming to improve efficiency by selecting a subset of features or data. Random subset selection, a simple and widely used method, involves randomly sampling data, often for cross-validation or bootstrap (Hastie, 2009). Importance-based selection focuses on high-value data points, while active learning targets samples that are expected to provide the most information, improving model learning (Quinlan, 1986). Filter methods rank features using statistical measures such as correlation or variance, selecting the top-ranked ones for modeling (Guyon & Elisseeff, 2003). Narendra & Fukunaga (1977) introduced a branch-and-bound algorithm for efficient feature selection, and Wei et al. (2015) proposed filtering active submodular selection (FASS), combining uncertainty sampling with submodular optimization. Yang et al. (2022) proposed dataset pruning, an optimization-based sample selection method that identifies the smallest subset of training data to minimize generalization gaps while significantly reducing training costs. Zhu (2016) proposed an efficient method to approximate the gradient of the objective function using a pilot estimate. The core idea is to compute the gradient information corresponding to each data point based on an initial parameter estimate (referred to as the “pilot estimate”) and identify data points with larger gradient values as more “important” samples for subsequent optimization. Despite these advancements, directly applying subset selection methods to BO often yields suboptimal results, necessitating further exploration to integrate subsampling effectively into BO frameworks.
3 Background
@@ 3.1 Bayesian Optimization and Gaussian Processes
BO aims to find the global optimum $x^{*}\in\mathcal{X}$ of an unknown reward function $f:\mathcal{X}\to\mathbb{R}$, over the $n$-dimensional input space $\mathcal{X}=[0,1]^{n}$. Throughout this paper, we consider minimization problems, i.e., we aim to find $x^{*}\in\mathcal{X}$ such that $f(x^{*})\leq f(x)$ for all $x\in\mathcal{X}$, approximating the performance of the optimal point $x^{*}=\arg\max_{x\in\mathcal{X}}f(x)$ as quickly as possible. GPs are one of the fundamental components in BO, providing a theoretical framework for modeling and prediction in a black-box function. In each round, a sample $x_{t}$ is selected based on the current GP’s posterior and acquisition function. The observed values $y_{t}$ and $x_{t}$ are then stored in the sample buffer, and the GP surrogate is updated according to these samples. This iterative process of sampling and updating continues until the optimization objects are achieved or the available budget is exhausted.
The key advantage of GPs lies in their nonparametric nature, allowing them to model complex functions without assuming a specific form. GPs are widely used for regression (Gaussian Process Regression (Schulz et al., 2018), GPR) and classification tasks due to their flexibility and ability to provide uncertainty estimates. Formally, a GP can be defined as: $f(\mathbf{x})\sim\mathcal{GP}(\mu(x),k(x,x^{\prime})))$, where $\mu(x)$ is the mean function, often assumed to be zero, and $k(x,x^{\prime})$ is the covariance function, defining the similarity between points $x$ and $x^{\prime}$. It should be noted that the algorithmic complexity of GP updates is $\mathcal{O}(n^{3})$, where $n$ is the number of observed samples. As the sample set grows, the computational resources required for these updates can become prohibitively expensive, especially in large-scale optimization problems.
@@ 3.2 Diversity-based Subset Selection
Due to limited computing resources, intelligently selecting samples instead of using all samples to fit a model is more efficient in problems with a large sample set. In continuing learning, this will overcome the catastrophic forgetting of previously seen data when faced with online data streams. Suppose that we have a model fitted on observed samples $\mathcal{D}\triangleq\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}\in\mathcal{X}$ and $y_{i}$ is the corresponding observation. In the context of subset selection, our objective is to ensure that each newly added sample contributes meaningfully to the optimization process; that is, the goal of sequential selecting the sample is to minimize the loss function $\ell(f(x_{n};\theta),y_{n})$, where $f(\cdot;\theta)$ denotes the model parameterized by $\theta$. Formally, this can be expressed as:
$\begin{split}\theta^{n}&=\arg\min_{\theta}\ell(f(x_{n};\theta),y_{n})\\ \text{s.t.}\quad&\ell(f(x_{i};\theta),y_{i})\leq\ell(f(x_{i};\theta^{n-1}),y_{i}),\,\\ &\forall i\in\{1,\dots,n-1\}.\end{split}$ (1)
The constraints ensure that when we select new samples for the sample subset, the loss of the new samples will not exceed that of the previous subset samples, thereby preserving the performance of the previously observed subset.
Let $g_{n}=\nabla_{\theta}\ell(f(x_{n};\theta),y_{n})$ be the gradient of loss for model parameters $\theta$ at time $n$. Following Aljundi et al. (2019), we rephrase the constraints involving the loss with respect to the gradients. Specifically, the constraint of (1) can be rewritten as $\langle\mathbf{g}_{n},\mathbf{g}_{i}\rangle\geq 0,\forall i\in\{1,\dots,n-1\}.$ This transformation simplifies the constraint by focusing on the inner product of the gradients, which are nonnegative such that the loss does not increase.
To solve (1), we consider the geometric properties of the gradients in the parameter space. Note that optimizing the solid angle subtended by the gradients is computationally expensive. According to the derivation in Aljundi et al. (2019), the sample selection problem is equivalent to maximizing the variance of the loss gradient direction of the samples in the fixed-size buffer. By maximizing the variance of the gradient directions, we ensure that the selected samples represent diverse regions of the parameter space, and therefore the buffer contains diverse samples, each contributing unique information to the optimization process. How to determine the buffer size will be detailed in Section 4.3. Problem (1) thus becomes a surrogate of selecting a subset $\mathcal{U}$ of the samples that maximize the diversity of their gradients:
@@
Here, $M$ denotes the number of samples saved in the buffer and $/\|\textbf{g}\|$ g ∥ is the normalized gradient vector. The reformulated problem (2) transforms the sample selection process from a sequential approach (adding samples to the subset one at a time) into a batch selection approach (samples are selected all at once).
@@ 4.1 Gradient Information with GP
In light of the pilot estimate-based gradient information acquisition method in Zhu (2016), we propose a new method for gradient information acquisition in GPs. In a GP model, given a set of samples $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ that follows a multivariate normal distribution with mean $\mu$ and covariance matrix $\mathbf{K}$, where $\mathbf{K}$ is constructed from a kernel function $k(\mathbf{x}_{i},\mathbf{x}_{j};\theta)$ and ${\theta}$ represents the hyperparameters, the probability density function of a multivariate Gaussian distribution is $p(\mathbf{y}|\mathbf{X},{\theta})=\frac{1}{(2\pi)^{n/2}|\mathbf{K}|^{1/2}}\exp\left(-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}\right)$.Taking logarithm on it and we derive the log-likelihood function:
$\begin{split}\log p(\mathbf{y}|\mathbf{X},{\theta})=&-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}\\ &-\frac{1}{2}\log|\mathbf{K}|-\frac{n}{2}\log(2\pi).\end{split}$ (3)
@@
The derivative of $\mathbf{y}$ directly measures how sensitive this log-likelihood is to each observation $y_{i}$. Intuitively, if changing $y_{i}$ significantly alters the value of (3), that sample has a large marginal contribution to the fit. Hence, in subset selection schemes, one can use these gradient magnitudes to gauge how important each sample is, potentially adjusting their weights or deciding which samples to retain in a subset.
To define the gradient for each sample, we first quantify each sample’s contribution to the log-likelihood. Since the second and third terms in (3), $-\frac{1}{2}\log|\mathbf{K}|$ and $-\frac{n}{2}\log(2\pi)$, do not depend on $\mathbf{y}$, both have no contribution to the gradient. Consequently, the gradient of the log-likelihood with respect to $\mathbf{y}$ is given by: $\frac{\partial\log p(\mathbf{y|X},{\theta})}{\partial\mathbf{y}}=-\mathbf{K}^{-1}(\mathbf{y}-\mu)$. Note that the $i$-th component, $-(\mathbf{K}^{-1}(\mathbf{y}-\mu))_{i}$, corresponds to the partial derivative of the log-likelihood with respect to $\mathbf{y}_{i}$. Thus, we define the gradient $g_{i}$ for each sample $\mathbf{y}_{i}$ as:
@@
This gradient calculation is computationally efficient as the value of $\mathbf{K}^{-1}$ in (4) is available while updating the GP. Furthermore, the complexity of the additional computational burden introduced by the gradient calculation is $O(n^{2})$, and it is negligible to the complexity of GP updates (which is $O(n^{3})$), especially when $n$ is large.

### v2

@@ Abstract
Bayesian optimization (BO) is an effective technique for black-box optimization. However, its applicability is typically limited to moderate-budget problems due to the cubic complexity in computing the Gaussian process (GP) surrogate model. In large-budget scenarios, directly employing the standard GP model faces significant challenges in computational time and resource requirements. In this paper, we propose a novel approach, gradient-based sample selection Bayesian Optimization (GSSBO), to enhance the computational efficiency of BO. The GP model is constructed on a selected set of samples instead of the whole dataset. These samples are selected by leveraging gradient information to remove redundancy while preserving diversity and representativeness. We provide a theoretical analysis of the gradient-based sample selection strategy and obtain explicit sublinear regret bounds for our proposed framework. Extensive experiments on synthetic and real-world tasks demonstrate that our approach significantly reduces the computational cost of GP fitting in BO while maintaining optimization performance comparable to baseline methods.
@@ 1 Introduction
Bayesian optimization (BO) is a derivative-free optimization method for black-box functions [1, 2, 3] that enjoys enormous success in many applications such as machine learning [4], structural design [5], robotics [6], additive manufacturing [7], inertial confinement fusion design  [8], etc. BO’s strength lies in its ability to represent the unknown objective function through a surrogate model and by optimizing an acquisition function [2, 3]. BO consists of a surrogate model, which provides a global prediction for the unknown objective function, and an acquisition function that serves as a criterion to determine the next sample to evaluate. In particular, the Gaussian process (GP) model is often preferred as the surrogate model due to its versatility and reliable uncertainty estimation. However, the GP model often suffers from large data sets, making it more suitable for small-budget scenarios [4]. To fit a GP model, the dominant complexity in computing the inversion of the covariance matrix is $\mathcal{O}(n^{3})$, where $n$ is the number of data samples. As the sample set grows, the computational burden increases substantially. This limitation poses a significant challenge for scaling BO to real-world problems with large sample sets.
@@
Despite the various approaches to improve the computational efficiency of BO, including parallel BO [5, 6, 9, 10], kernel approximation [11, 12, 13, 14] and sparse GP [15, 16, 17], the computational overhead remains a burden in practice [18]. Kernel approximation methods typically involve simplifying or approximating the kernel matrix, which can degrade accuracy and lead to suboptimal performance if the approximation is too coarse. Sparse GP methods, while reducing computational complexity, introduce additional complexity through corrective terms to maintain approximation accuracy, and existing iterative implementations in Sparse GP [17] frequently involve adding new samples while removing previous ones, potentially causing inefficiencies or suboptimal sample selection in complex optimization landscapes.
During the iterative search process of BO, some samples can become redundant and contribute little to the additional information gain. Such samples collected in earlier stages thus diminish in importance as the process evolves. For instance, excessive searching around identified minima becomes redundant once the optimal value has been determined, as these samples cease to offer meaningful insights for further optimization. To efficiently fit a GP, it is essential to focus on samples that provide the most informative contributions. As shown in Figure 1, carefully selected samples can effectively fit a GP. Despite the reduced number of samples, the GP still captures the key trends and features of the true function while maintaining reasonable uncertainty bounds. In this paper, we propose incorporating the gradient-based sample selection technique into the BO framework to enhance its scalability and effectiveness in large-budget scenarios. This technique was originally proposed for continual learning with online data stream [19]. The previous data are selectively sampled and stored in a replay buffer to prevent catastrophic forgetting and enhance model fitting. By using gradient information to gauge the value of each sample, one can more judiciously decide which samples are most essential for building a subset, and maintain the most representative subset of BO samples. This subset is then used to fit the GP model, accelerating the BO process while ensuring efficient and effective GP fitting. We summarize our main contributions as follows:
@@
  • Theoretical analysis. We provide a rigorous theoretical analysis of the regret bound for the GSSBO. Theoretical results show that the regret bound of our proposed algorithm is similar to that of the standard GP-UCB algorithm.
  • Empirical validations. We conduct comprehensive numerical experiments, including synthetic and real-world test problems, to demonstrate that compared to baseline methods. The proposed algorithm achieves comparable performance and significantly reduces computational costs. These results verify the benefit of using gradient information to select a representative subset of samples.
@@ 2 Related Works
BO with Resource Challenges. In practical applications, BO faces numerous challenges, including high evaluation costs, input-switching costs, resource constraints, and high-dimensional search spaces. Researchers have proposed a variety of methods to address these issues. For instance, parallel BO employs batch sampling to improve efficiency in large-scale or highly concurrent scenarios [5, 6, 9, 10]. Kernel approximation methods, such as random Fourier features, map kernels onto lower-dimensional feature spaces, thus accelerating kernel-based approaches [20, 11]. Multi-fidelity BO leverages coarse simulations with a limited number of high-fidelity evaluations to reduce the overall cost [21]. For high-dimensional tasks, random embeddings or active subspaces help reduce the search dimensionality [22, 23]. Meanwhile, sparse GP significantly reduces computational complexity by introducing inducing points [15, 16, 17, 24]. However, these approaches face limitations in practical scenarios and usually sacrifice performance for scaling. Calandriello et al. [25] scaling GP optimization by repeatedly evaluating each selected point until its posterior uncertainty falls below a preset threshold, thus limiting the number of datasets. However, the dataset still grows with time, and the algorithm’s dependence on its initial sample set means that low-value points selected early on remain permanently in the model, potentially inflating computational overhead.
@@
BO with Gradient Information. The availability of derivative information can significantly simplify optimization problems. Ahmed et al. [26] highlights the potential of incorporating gradient information into BO methods and advocates for its integration into optimization frameworks. Wu et al. [27] introduced the parallel knowledge gradient method for batch BO, achieving faster convergence to global optima. Then they introduce d-KG [28] as a new acquisition function, which systematically introduces gradient information into the BO and proves that using gradient information can strictly improve the information value. Rana et al. [29] incorporated GP priors to enable gradient-based local optimization. Chen et al. [30] proposed a unified particle-optimization framework using Wasserstein gradient flows for scalable Bayesian sampling. Bilal et al. [31] demonstrated that BO with gradient-boosted regression trees performed well in cloud configuration tasks. Tamiya et al. [32] developed stochastic gradient line BO (SGLBO) for noise-robust quantum circuit optimization. Penubothula et al. [33] funded local critical points by querying where the predicted gradient is zero. Zhang et al. [34] introduced BO of gradient trajectory (BOGAT) for efficient imaging optimization. Makrygiorgos et al. [35] integrate exact gradient observations into the Bayesian neural network surrogate’s training loss. Although these methods leverage gradient information to improve optimization efficiency and performance, they mainly focus on refining the GP model or acquisition functions.
@@
Subset Selection. Subset selection is a key task in fields such as regression, classification, and model selection, aiming to improve efficiency by selecting a subset of features or data. Random subset selection, a simple and widely used method, involves randomly sampling data, often for cross-validation or bootstrap [36]. Importance-based selection focuses on high-value data points, while active learning targets samples that are expected to provide the most information, improving model learning [37]. Filter methods rank features using statistical measures such as correlation or variance, selecting the top-ranked ones for modeling [38]. Narendra et al. [39] introduced a branch-and-bound algorithm for efficient feature selection. Yang et al. [40] proposed dataset pruning, an optimization-based sample selection method that identifies the smallest subset of training data to reduce training costs. Ash et al. [41] employs the k-means++ algorithm in the gradient space for diversity sampling in active learning. Oglic [42]first maps each data point into the reproducing kernel Hilbert space (RKHS), then uses a max–min coverage strategy in the RKHS to sequentially sample $K$ landmarks, which are employed to construct the Nyström low-rank approximation. This method is dependent on the quality of the selected landmarks. Hayakawa [43] provides tighter expected error bounds under a continuous measure for the same underlying idea. However, computing the Mercer decomposition in high dimensions incurs substantial computational cost and suffers from severe error degradation. Because both of these methods perform their approximations in RKHS at high computational expense, they are most suitable for offline, batch-mode resampling. Zhu et al. [44] proposed a “pilot estimate” to approximate the gradient of the objective function. The core idea is to compute the gradient information corresponding to each data point based on an initial parameter estimate and identify data points with larger gradient values as more “important” samples for subsequent optimization. Despite these advancements, directly applying subset selection methods to BO often yields suboptimal results, necessitating further exploration to integrate subsampling effectively into BO frameworks.
3 Preliminaries
@@ 3.1 Bayesian Optimization and Gaussian Processes
BO aims to find the global optimum $x^{*}\in\mathcal{X}$ of an unknown reward function $f:\mathcal{X}\to\mathbb{R}$, over the $n$-dimensional input space $\mathcal{X}=[0,1]^{n}$. Throughout this paper, we consider maximization problems, i.e., we aim to find $x^{*}\in\mathcal{X}$ such that $f(x^{*})\leq f(x)$ for all $x\in\mathcal{X}$, get the optimal point $x^{*}=\arg\max_{x\in\mathcal{X}}f(x)$ as quickly as possible. GPs are one of the fundamental components in BO, providing a theoretical framework for modeling and prediction in a black-box function. In each round, a sample $x_{t}$ is selected based on the current GP’s posterior and acquisition function. The observed values $y_{t}$ and $x_{t}$ are then stored in the sample buffer, and the GP surrogate is updated according to these samples. This iterative process of sampling and updating continues until the optimization objectives are achieved or the available budget is exhausted. The key advantage of GPs lies in their nonparametric nature, allowing them to model complex functions without assuming a specific form. GPs are widely used for regression (Gaussian Process Regression [45], GPR) and classification tasks due to their flexibility and ability to provide uncertainty estimates. Formally, a GP can be defined as: $f(\mathbf{x})\sim\mathcal{GP}(\mu(x),k(x,x^{\prime})))$, where $\mu(x)$ is the mean function, often assumed to be zero, and $k(x,x^{\prime})$ is the covariance function, defining the similarity between points $x$ and $x^{\prime}$. It should be noted that the algorithmic complexity of GP updates is $\mathcal{O}(n^{3})$, where $n$ is the number of observed samples. As the sample set grows, the computational resources required for these updates can become prohibitively expensive, especially in large-scale optimization problems.
@@ 3.2 Diversity-based Subset Selection
Due to limited computing resources, properly selecting samples instead of using all samples to fit a model is more efficient in problems with a large sample set. In continual learning, this helps overcome the catastrophic forgetting of previously seen data when faced with online data streams. Suppose that we have a model fitted on observed samples $\mathcal{D}\triangleq\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}\in\mathcal{X}$ and $y_{i}$ is the corresponding observation. In the context of subset selection, our objective is to ensure that each newly added sample contributes meaningfully to the optimization process. That is, a constraint ensures that when we select new samples for the sample subset, the performance of the model after the new samples are added will not be worse than the performance of the previous subset samples. Let $\mathbf{g}_{n}$ be the gradient of the sample at time $n$. Following [19], we rephrase the constraints with respect to the gradients. Specifically, the constraint can be rewritten as $\langle\mathbf{g}_{n},\mathbf{g}_{i}\rangle\geq 0,\forall i\in\{1,\dots,n-1\}.$ This transformation simplifies the constraint by focusing on the inner product of the gradients, which are nonnegative, such that there will be no performance degradation. To solve the constraint, we can use the geometric properties of the gradients in the parameter space. Note that optimizing the solid angle subtended by the gradients is computationally expensive. According to the derivation in [19], the sample selection problem is equivalent to maximizing the variance of the gradient direction of the samples in the fixed-size buffer. By maximizing the variance of the gradient directions, we ensure that the selected samples represent diverse regions of the parameter space, and therefore the buffer contains diverse samples, each contributing unique information to the optimization process. How to determine the buffer size will be detailed in Section 4.3. The previous problem thus becomes a surrogate for selecting a subset $\mathcal{U}$ of the samples that maximizes the diversity of their gradients:
@@
Here, $M$ denotes the buffer size and $/\|\textbf{g}\|$ g ∥ is the normalized gradient vector. The larger the value of this formula, the more dispersed the gradient direction of the selected samples is (the higher the diversity is). The reformulated problem (1) transforms the sample selection process from a sequential approach (adding samples to the subset one at a time) into a batch selection approach (samples are selected all at once). This empirical surrogate objective is agnostic to how gradient information is computed, making it straightforward to integrate into subset-based methods.
@@ 4.1 Gradient Information with GP
In the previous section, we introduced how to use the gradient to select a diverse subset of samples. However, within the Bayesian optimization framework using Gaussian processes, we are more concerned with the impact on the GP posterior of the information carried by the observation $y_{i}$. Therefore, in light of the pilot estimate-based gradient information acquisition method in [44], we propose a new method for gradient information acquisition in GPs: $g_{i}\;=\;\frac{\partial}{\partial y_{i}}\,\log p(\mathbf{y}\mid\mathbf{X},\theta).$ We focus on the gradient of output $\mathbf{y}$ rather than input $\mathbf{X}$ as the derivative of $\mathbf{y}$ is simpler and directly reflects each sample’s contribution. In a GP model, given a set of samples $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ that follows a multivariate normal distribution with mean $\mu$ and covariance matrix $\mathbf{K}$, where $\mathbf{K}$ is constructed from a kernel function $k(\mathbf{x}_{i},\mathbf{x}_{j};\theta)$ and ${\theta}$ represents the hyperparameters, the probability density function of a multivariate Gaussian distribution is $p(\mathbf{y}|\mathbf{X},{\theta})=\frac{1}{(2\pi)^{n/2}|\mathbf{K}|^{1/2}}\exp\left(-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}\right)$. Taking the logarithm of it, we derive the log-likelihood function:
$\begin{split}\log p(\mathbf{y}|\mathbf{X},{\theta})=-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}-\frac{1}{2}\log|\mathbf{K}|-\frac{n}{2}\log(2\pi).\end{split}$ (2)
@@
The derivative of log likelihood with respect to $\mathbf{y}$ directly measures how sensitive this log-likelihood is to each observation $y_{i}$. Intuitively, if changing $y_{i}$ significantly alters the value of (2), that sample has a large marginal contribution to the fit. Hence, in subset selection schemes, one can use these gradient magnitudes to gauge how important each sample is, potentially adjusting their weights or deciding which samples to retain in a subset. To define the gradient for each sample, we first quantify each sample’s contribution to the log-likelihood. Since the second and third terms in (2), $-\frac{1}{2}\log|\mathbf{K}|$ and $-\frac{n}{2}\log(2\pi)$, do not depend on $\mathbf{y}$, both have no contribution to the gradient. Consequently, the gradient of the log-likelihood with respect to $\mathbf{y}$ is given by: $\frac{\partial\log p(\mathbf{y|X},{\theta})}{\partial\mathbf{y}}=-\mathbf{K}^{-1}(\mathbf{y}-\mu)$. Note that the $i$-th component, $-(\mathbf{K}^{-1}(\mathbf{y}-\mu))_{i}$, corresponds to the partial derivative of the log-likelihood with respect to $\mathbf{y}_{i}$. Thus, we define the gradient $g_{i}$ for each sample $\mathbf{y}_{i}$ as:
@@
After obtaining the gradient information of each sample, we can use the maximization of the variance of the angle between the two gradients as an alternative goal (i.e. maintaining the diversity of gradient directions). This gradient calculation is computationally efficient as the value of $\mathbf{K}^{-1}$ in (3) is available while updating the GP. Furthermore, the complexity of the additional computational burden introduced by the gradient calculation is $O(n^{2})$, and it is negligible compared to the complexity of GP updates (which is $O(n^{3})$), especially when $n$ is large.